# Continuous Diffusion for Mixed-Type Tabular Data

### v2

@@ Appendix D Derivation of the Functional Timewarping Form
Since higher noise levels, $\sigma$, imply a lower signal-to-noise ratio and, in turn, a larger loss, $\ell$, we know that the loss must be a monotonically increasing and S-shaped function of the noise level. Additionally, the function has to be easy to invert and differentiate. We incorporate this prior information in the functional timewarping form of $F:\sigma\mapsto\ell$. A convenient choice is the cdf of the logistic distribution:
@@
Appendix E Datasets
Our selected benchmark datasets are highly diverse, particularly in the number of categories for a categorical feature. Almost all datasets are licensed under creative commons that permits their usage in our research. Only the nmes dataset license is unknown but the data is publicly accessible on the author’s website.
For the diabetes and covertype datasets, we transform the original multi-class classification problem into a binary classification task for easier comparison. For the covertype data, the task is converted into predicting whether a forest of type 2 is present in a given 30 $\times$ 30 meter area. In the diabetes data, we convert the task by predicting whether a patient was readmitted to a hospital.
@@ Appendix F Baseline Models
Below, we give a brief description of the used generative baseline models (including code sources).
@@
(Lee et al., 2023) – a diffusion model trained using an additional contrastive loss, and which factorizes the joint distribution of mixed-type tabular data into a distribution for continuous data conditional on categorical features and a distribution for categorical data conditional on continuous features. Similar to TabDDPM, the authors utilize the multinomial diffusion framework (Hoogeboom et al., 2021) to model categorical data. An implementation is available at https://github.com/ChaejeongLee/CoDi under an unknown license.
@@ Appendix G Implementation Details
Each of the selected benchmark models requires a rather different, more specialized neural network architecture. Imposing the same architecture across models is therefore not possible. The same inability holds for the comparison of CDTD to other diffusion-based models: Our model is the first to use a continuous noise distribution on both continuous and categorical features, and therefore the alignment of important design choices, like the noise schedule, across models is not possible. In particular, the Markov transition matrices based forward process of the multinomial diffusion framework (Hoogeboom et al., 2021) in TabDDPM and CoDi does not translate to our setting.
To ensure a fair comparison in terms of sampling steps, we set the steps for CDTD, TabDDPM and CoDi to $\max(200,\text{default})$. We therefore increase the default number of sampling steps for CoDi (from 50 steps) and TabDDPM (from 100 steps for classification datasets). For TabDDPM and regression datasets, we use the suggested default of 1000 sampling steps.
@@
We adjust each architecture to a total of ±$\pm 3$ million trainable parameters on the adult dataset to improve the comparability further (see Table 4) and use the same architectures for all considered datasets. Note that the total number of parameters may vary slightly across datasets due to different number of features and categories affecting the onehot encoding but is still comparable across models. We also align the embedding/bottleneck dimensions for CTGAN, TVAE, TabDDPM and CDTD to 256 to accommodate more complex datasets. If applicable, all models are trained for 30k steps on a single RTX 4090 instance, using PyTorch version 2.2.2.
@@
(Xu et al., 2019): We follow the popular implementation in the Synthetic Data Vault package (see https://github.com/sdv-dev/CTGAN). For this model to work, the batch size must be divisible by 10. Therefore, we adjust the batch size if necessary. We use a 258-dimensional embedding (instead of the default embedding dimension of 128) to better align the CTGAN architecture with TVAE, TabDDPM and CDCD.
@@
(Xu et al., 2019): We again follow the implementation in the Synthetic Data Vault. We use a 258-dimensional embedding to better align the architecture with CTGAN, TabDDPM and CDCD.
@@
(Kotelnikov et al., 2023): There are no general default hyperparameters provided. Hence, we mostly adapt the papers’ tuned hyperparameters for the adult dataset (one of the few used datasets that includes both continuous and categorical features). However, we decrease the learning rate from 0.002 to 0.001, since most of the tuned models in the paper used learning rates around 0.001. For regression task datasets, we use 1000 sampling steps in accordance with the author’s settings. For classification task datasets, we use 200 sampling steps (instead of the default 100 steps), to better align the model with CoDi and CDCD. Note also that for classification task datasets, TabDDPM models the conditional distribution $p(x|y)$, instead of the unconditional distribution $p(x)$ which is modeled for regression tasks. We adjust the dimension of the bottleneck to 256 (instead of the default 128) to also accommodate also larger datasets and align the model with CTGAN, TVAE,and CDCD.
@@
(ours): To ensure comparability in particular to TabDDPM and CoDi, we use the same neural network architecture as TabDDPM. We only change the input layers to accommodate our embedding-based framework. In the input layer, we vectorize all embedded categorical features and concatenate them with the scalar valued continuous features. The adjusted output layer ensures that we predict a single value for each continuous features and set of class-specific probabilities for each categorical feature. Since our use of embeddings introduces additional parameters, we scale the hidden layers slightly down relative to the TabDDPM to ensure approximately 3 million trainable parameters (instead of 808 neurons per layer we use 806) on the adult dataset. More details on the CDTD implementation are given in Appendix J.
@@ Appendix H Tuning of the Detection Model
We use a catboost model (Prokhorenkova et al., 2018) to test whether real and generated samples can be distinguished. We generate the same number of fake observations for each of the real train, validation and test sets. We cap the maximum size of the real data subsets to 25 000, and subsample them if necessary, to limit the computational load. We then combine real and fake observations per set to $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$, and $\mathcal{D}_{\text{test}}^{\text{detect}}$, respectively. The catboost model is trained on $\mathcal{D}_{\text{train}}^{\text{detect}}$ with the task of predicting whether an observation is real or fake. We tune the catboost model with optuna and for 50 trials to maximize the accuracy on $\mathcal{D}_{\text{valid}}^{\text{detect}}$. The catboost hyperparameter search space is given in Table 5. Afterwards, we repeat the sampling process and the creation of $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$ and $\mathcal{D}_{\text{test}}^{\text{detect}}$ for five different seeds. Each time, the model is trained on $\mathcal{D}_{\text{train}}^{\text{detect}}$ with the previously tuned hyperparameters, and evaluated on $\mathcal{D}_{\text{test}}^{\text{detect}}$. The average test set accuracy over the five seeds yields the estimated detection score.
@@ Appendix I Machine Learning Efficiency Models
To enable a fair comparison to the other methods and in particular TabDDPM, the CDTD score model utilizes the exact same architecture and optimizer as Kotelnikov et al. (2023). An overview of the score model is given in Figure 4: First, the noisy data, including the noisy scalars for continuous features and the noisy embeddings for categorical features, are projected onto a 256-dimensional space. Similarly, timestep $t$ and possibly conditioning information $y$ are embedded in the same space. All 256-dimensional vectors are then added and the results is processed by a set of five fully-connected linear layers with ReLU activation functions. Lastly, a linear projection maps the output of the fully-connected layers to the required output dimensions, which depend on the number of features and number of categories per feature.
@@
The only major difference to the TabDDPM setup are the inputs, as we need to embed the categorical features in Euclidean space. The output dimensions are the same, as we need to predict a single scalar for each $x_{\text{cont},i}$, and $\lvert C_{j}\rvert$ values for each $x_{\text{cat},j}$, with $C_{j}$ the set of categories of feature $j$. We change the initialization of the output layer as described in Appendix B: To handle our inputs, we embed the categorical features in 16-dimensional space and add a feature-specific bias of the same dimension, which captures feature-specific information common to all categories and is initialized to zero. We L2-normalize each embedding to prevent a degenerate embedding space in which embeddings are pushed further and further apart (see Dieleman et al., 2022). Also, Dieleman et al. (2022) argue that the standard deviation of the Normal distribution used to initialize the embeddings, denoted by $\sigma_{\text{cat emb}}$, is an important hyperparameter. In this paper, we set $\sigma_{\text{cat emb}}=0.001$ for all datasets and have not seen detrimental effects. Tuning this hyperparameter in practice could improve the results further.
@@
Since we utilize embeddings, we have to scale the neurons per layer slightly down in the stack of the five fully-connected layers (from 808 for TabDDPM to 806). Also, since TabDDPM samples discrete steps from $[0,T]$, with $T\gg 1$, we adjust our timesteps $t\in[0,1]$ up by 1000. We use the same optimizer (Adam), learning rate (0.001), learning rate decay (linear), EMA decay (0.999), and training steps (30000). However, since we work with embeddings we add a linear warmup schedule over the first 100 steps.
@@
Instead of using the vanilla uniform (time)step sampling as the TabDDPM, the CDTD model uses antithetic sampling Dieleman et al. (2022); Kingma et al. (2022). The timesteps are still uniformly distributed but spread out more evenly over the domain, which benefits the training of the adaptive noise schedules. For generation, we use an Euler sampler with 200 steps to minimize the discretization error.
Appendix K Examples of Learned Noise Schedules
@@
Below we show the learned noise schedules for the smallest dataset (churn) and the largest dataset (acsincome). We also illustrate how well single, per type and per feature schedules fit the respective losses for each of the datasets.
@@
Appendix L Qualitative Comparisons
@@
Figure 10: Qualitative comparison of some univariate distributions for adult, bank, churn, nmes data.
Figure 11: Qualitative comparison of some univariate distributions for beijing, default, lending, news data. Note that CoDi is excluded from lending as it is prohibitively expensive to train on this dataset.
Appendix M Visualizations of Captured Correlations
Figure 12: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the acsincome dataset. TabDDPM generates NaNs for this dataset and is therefore excluded. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
@@
Figure 17: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the covertype dataset. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
@@
Appendix N Detailed Results

### v3

@@ Appendix D Derivation of the Functional Timewarping Form
Since higher noise levels, $\sigma$, imply a lower signal-to-noise ratio, and in turn a larger loss, $\ell$, we know that the loss must be a monotonically increasing and S-shaped function of the noise level. Additionally, the function has to be easy to invert and differentiate. We incorporate this prior information in the functional timewarping form of $F:\sigma\mapsto\ell$. A convenient choice is the cdf of the logistic distribution:
@@
Appendix E Benchmark Datasets
Our selected benchmark datasets are highly diverse, particularly in the number of categories for categorical features (see Table 3). For the diabetes and covertype datasets, we transform the original multi-class classification problem into a binary classification task for ease of presentation. For the covertype data, the task is converted into predicting whether a forest of type 2 is present in a given 30 $\times$ 30 meter area. In the diabetes data, we convert the task by predicting whether a patient was readmitted to a hospital. All datasets are publicly accessible and (except nmes) licensed under creative commons.
@@ Appendix F Baseline Models
Below, we give a brief description of our selected generative baseline models (including code sources).
SMOTE
(Chawla et al., 2002) – a technique (not a generative model) typically used to oversample minority classes based on interpolation between ground-truth observations. We use SMOTENC for mixed-type data from the scikit-learn package and mostly adapt the code from the TabDDPM repository (Kotelnikov et al., 2023). For sampling, we utilize 16 CPU cores.
@@
(Lee et al., 2023) – a diffusion model trained with an additional contrastive loss, and which factorizes the joint distribution of mixed-type tabular data into a distribution for continuous data conditional on categorical features and a distribution for categorical data conditional on continuous features. Similarly, the authors utilize the multinomial diffusion framework (Hoogeboom et al., 2021) to model categorical data. An implementation is available at https://github.com/ChaejeongLee/CoDi under an unknown license.
TabSyn
(Zhang et al., 2024) – a diffusion-based model that first learns a tranformer-based VAE to map mixed-type data to a continuous latent space. Then, the diffusion model is trained on that latent space. We use the official code available at https://github.com/amazon-science/tabsyn under the Apache 2.0 license.
@@ Appendix G Implementation Details
Each of the selected benchmark models requires a rather different, more specialized neural network architecture. Imposing the same architecture across models is therefore not possible. The same inability holds for the comparison of CDTD to other diffusion-based models: Our model is the first to use a continuous noise distribution on both continuous and categorical features, and therefore the alignment of important design choices, like the noise schedule, across models is not possible. In particular, the forward process of the multinomial diffusion framework (Hoogeboom et al., 2021) used in TabDDPM and CoDi, which is based on Markov transition matrices, does not translate to our setting.
To ensure a fair comparison in terms of sampling steps, we set the steps for CDTD, TabDDPM, CoDi and TabSyn to $\max(200,\text{default})$. We therefore increase the default number of sampling steps for CoDi and TabSyn (from 50 steps) and TabDDPM (from 100 steps for classification datasets). For TabDDPM and regression datasets, we use the suggested default of 1000 sampling steps.
@@
We adjust each architecture to a total of ±$\pm 3$ million trainable parameters on the adult dataset to improve the comparability further (see Table 4) and use the same architectures for all considered datasets. Note that the total number of parameters may vary slightly across datasets due to different number of features and categories affecting the onehot encoding but is still comparable across models. We also align the embedding/bottleneck dimensions for CTGAN, TVAE, TabDDPM, TabSyn and CDTD to 256. To align TabDDPM, TabSyn and CDTD further, we use the TabDDPM architecture for all models, with appropriate adjustments for different input types and dimensions. If applicable, all models are trained for 30k steps on a single RTX 4090 instance, using PyTorch version 2.2.2.
@@
SMOTE
(Chawla et al., 2002): We use the default hyperparameters suggested for the SMOTENC scikit-learn implementation.
@@
(Xu et al., 2019): We follow the popular implementation in the Synthetic Data Vault package (see https://github.com/sdv-dev/CTGAN). For this model to work, the batch size must be divisible by 10. Therefore, we adjust the batch size if necessary. We use a 256-dimensional embedding (instead of the default embedding dimension of 128) to better align the CTGAN architecture with TVAE, TabDDPM, TabSyn and CDTD.
@@
(Xu et al., 2019): We again follow the implementation in the Synthetic Data Vault. We use a 256-dimensional embedding to better align the architecture with CTGAN, TabDDPM, TabSyn and CDTD.
@@
(Kotelnikov et al., 2023): There are no general default hyperparameters provided. Hence, we mostly adapt the papers’ tuned hyperparameters for the adult dataset (one of the few used datasets that includes both continuous and categorical features). However, we decrease the learning rate from 0.002 to 0.001, since most of the tuned models in the paper used learning rates around 0.001. For regression task datasets, we use 1000 sampling steps in accordance with the author’s settings. For classification task datasets, we use 200 sampling steps (instead of the default 100 steps), to better align the model with CoDi and CDCD. Note also that for classification task datasets, TabDDPM models the conditional distribution $p(x|y)$, instead of the unconditional distribution $p(x)$ which is modeled for regression tasks. We adjust the dimension of the bottleneck to 256 (instead of the default 128) to also accommodate also larger datasets and align the model with CTGAN, TVAE,and CDTD.
@@
TabSyn
(Zhang et al., 2024): We use the default hyperparameters as suggested by the authors. The training steps that go towards training the VAE and the denoising network follow the proportions given in the official code (see https://github.com/amazon-science/tabsyn). To improve comparability to TabDDPM, CoDi and CDTD, we use the same neural network architecture as TabDDPM, which only differs slightly from the original architecture. We leave the VAE untouched.
@@
(ours): To ensure comparability in particular to TabDDPM, CoDi and TabSyn, we use the same neural network architecture as TabDDPM. We only change the input layers to accommodate our embedding-based framework. In the input layer, we vectorize all embedded categorical features and concatenate them with the scalar valued continuous features. The adjusted output layer ensures that we predict a single value for each continuous features and set of class-specific probabilities for each categorical feature. Since our use of embeddings introduces additional parameters, we scale the hidden layers slightly down relative to the TabDDPM to ensure approximately 3 million trainable parameters (instead of 808 neurons per layer we use 806) on the adult dataset. More details on the CDTD implementation are given in Appendix J.
@@ Appendix H Tuning of the Detection Model
We use a catboost model (Prokhorenkova et al., 2018) to test whether real and generated samples can be distinguished. We generate the same number of fake observations for each of the real train, validation and test sets. We cap the maximum size of the real data subsets to 25 000, and subsample them if necessary, to limit the computational load. Per set, we combine real and fake observations to $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$, and $\mathcal{D}_{\text{test}}^{\text{detect}}$, respectively. The catboost model is trained on $\mathcal{D}_{\text{train}}^{\text{detect}}$ with the task of predicting whether an observation is real or fake. We tune the catboost model with optuna and for 50 trials to maximize the accuracy on $\mathcal{D}_{\text{valid}}^{\text{detect}}$. The catboost hyperparameter search space is given in Table 5. Afterwards, we repeat the sampling process and the creation of $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$ and $\mathcal{D}_{\text{test}}^{\text{detect}}$ for five different seeds. Each time, the model is trained on $\mathcal{D}_{\text{train}}^{\text{detect}}$ with the previously tuned hyperparameters, and evaluated on $\mathcal{D}_{\text{test}}^{\text{detect}}$. The average test set accuracy over the five seeds yields the estimated detection score.
@@ Appendix I Machine Learning Efficiency Models
To enable a fair comparison to the other methods, and to TabDDPM and TabSyn in particular, the CDTD score model utilizes the exact same architecture and optimizer as Kotelnikov et al. (2023), which was also adapted by TabSyn (Zhang et al., 2024). An overview of the score model is provided in Figure 4: First, the noisy data, i.e., the noisy scalars for continuous features and the noisy embeddings for categorical features, are projected onto a 256-dimensional space. Similarly, timestep $t$ and possibly conditioning information $y$ are embedded in the same space. Then, all 256-dimensional vectors are added and the results is processed by a set of five fully-connected linear layers with ReLU activation functions. Lastly, a linear projection maps the output of the fully-connected layers to the required output dimensions, which depend on the number of features and number of categories per feature.
@@
The only major difference to the TabDDPM setup are the inputs, as we need to embed the categorical features in Euclidean space. The output dimensions are the same, as we need to predict a single scalar for each $x_{\text{cont},i}$, and $\lvert C_{j}\rvert$ values for each $x_{\text{cat},j}$, with $C_{j}$ the set of categories of feature $j$. We change the initialization of the output layer as described in Appendix B: To handle our inputs, we embed the categorical features in 16-dimensional space and add a feature-specific bias of the same dimension, which captures feature-specific information common to all categories and is initialized to zero. We L2-normalize each embedding to prevent a degenerate embedding space in which embeddings are pushed further and further apart (see Dieleman et al., 2022). Also, Dieleman et al. (2022) argue that the standard deviation of the Normal distribution used to initialize the embeddings, denoted by $\sigma_{\text{init}}$, is an important hyperparameter. In this paper, we set $\sigma_{\text{init}}=0.001$ for all datasets and have not seen detrimental effects. Table 7 indicates that CDTD is not sensitive to the choice of $\sigma_{\text{init}}$.
@@
Since we utilize embeddings, we have to scale the neurons per layer slightly down in the stack of the five fully-connected layers (from 808 for TabDDPM to 806). Also, since TabDDPM samples discrete steps from $[0,T]$, with $T\gg 1$, we scale our timesteps $t\in[0,1]$ up by 1000. We use the same optimizer (Adam), learning rate (0.001), learning rate decay (linear), EMA decay (0.999), and training steps (30000). However, since we work with embeddings we add a linear warmup schedule over the first 100 steps.
@@
Instead of using the vanilla uniform (time)step sampling as the TabDDPM, the CDTD model uses antithetic sampling (Dieleman et al., 2022; Kingma et al., 2022). The timesteps are still uniformly distributed but spread out more evenly over the domain, which benefits the training of the adaptive noise schedules. For generation, we use an Euler sampler with 200 steps to minimize the discretization error.
Appendix K CDTD Sampling
To sample from our learned distribution, we need to run the reverse process of the probability flow ODE (Equation 2). For example, for two different features $x_{1}$ and $x_{2}$, we deconstruct the ODE as:
$d{\mathbf{x}}$ $=-\frac{1}{2}{\mathbf{G}}(t){\mathbf{G}}(t)^{\mathsf{T}}\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}})dt$
$=-\begin{bmatrix}\dot{\sigma_{1}}(t)\sigma_{1}(t)&\\ &\dot{\sigma_{2}}(t)\sigma_{2}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)^{2}}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)^{2}}\end{bmatrix}dt$
$=-\begin{bmatrix}\dot{\sigma_{1}}(t)&\\ &\dot{\sigma_{2}}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)}\end{bmatrix}dt$
In practice, we use an Euler sampler with 200 discrete timesteps $\Delta t=t_{i+1}-t_{i}<0$. The timesteps are generated as a linearly spaced grid on $[0,1]$ and transformed afterwards into noise levels $\sigma_{k}(t)$ via the described timewarping procedure. For the discretized and simplified the ODE above, this yields
${\mathbf{x}}_{i+1}$ $={\mathbf{x}}_{i}-\begin{bmatrix}\frac{\Delta\sigma_{1}(t)}{\Delta t}&\\ &\frac{\Delta\sigma_{2}(t)}{\Delta t}\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)}\end{bmatrix}\Delta t={\mathbf{x}}_{i}+\begin{bmatrix}\frac{x_{1}-\hat{x_{1}}}{\sigma_{1}(t)}\\ \frac{x_{2}-\hat{x_{2}}}{\sigma_{2}(t)}\end{bmatrix}\odot\begin{bmatrix}\Delta\sigma_{1}(t)\\ \Delta\sigma_{2}(t)\end{bmatrix}.$
where $\odot$ denotes the element-wise product. Hence, we are effectively taking feature-specific steps of length $\Delta\sigma_{k}(t)$. The adaptive noise schedules (timewarping) therefore not only affect the training process, but also focus most work in the reverse process on the noise levels that matter most for sample quality (i.e., where $\Delta\sigma_{i}(t)$ is small).
We use finite differences to approximate $\dot{\sigma}_{i}$, instead of the available, analytical variant, since $\frac{d\sigma_{k}(t)}{dt}\rightarrow\infty$ as $t\rightarrow 1$. The step $\Delta t$ would therefore be required to decrease as $t\rightarrow 1$ to ensure $\Delta t\approx dt$ holds. For a large number of steps, this assumption does not hold in practice, and for $\frac{d\sigma_{k}(t)}{dt}$ the update of ${\mathbf{x}}$ overshoots the target drastically. Intuitively, $\sigma_{k}(t)$ becomes too steep near the terminal timestep $t=1$ such that the step size can not sufficiently compensate for the slope increase to turn $\frac{d\sigma_{k}(t)}{dt}$ into a good approximation of the actual change in $\sigma_{k}(t)$. Moreover, the analytical solution would approximate $d\sigma_{k}(t)=\dot{\sigma}_{k}(t)dt$, i.e., the change in the noise level caused by a change in $t$. Since we know exactly where $\sigma_{k}(t)$ will end up when changing $t$, we are better off using that exact value and let $d\sigma_{k}(t)=\Delta\sigma_{k}(t)$. Table 6 shows that the gains in sample quality are marginal to non-existent after more than 500 sampling steps.
Table 6: Performance sensitivity of CDTD (per type) to increasing number of sampling steps. Each metric is averaged over five seeds. As a robust measure, we report the median over the ablation study datasets acsincome, adult, beijing and churn .
Steps RMSE F1 AUC L2 distance of corr. Detection score JSD WD DCR
200 (default) 0.033 0.015 0.004 0.127 0.560 0.013 0.003 0.372
500 0.028 0.018 0.005 0.130 0.565 0.012 0.003 0.372
1000 0.028 0.018 0.005 0.129 0.560 0.012 0.003 0.373
1500 0.028 0.018 0.005 0.129 0.561 0.012 0.003 0.374
Appendix L Sensitivity to Important Hyperparameters
The training and sampling processes of CDTD are affected by various novel hyperparameters. Generally, a per-type noise schedule works best as we show in our main results in Table 1 for a diverse set of benchmark datasets. Here, we examine the sensitivity of CDTD to two additional important hyperparameters: (1) the standard deviation of the noise used to initialize the embeddings (and therefore specific to score interpolation), $\sigma_{\text{init}}$, and (2) the weight of the low noise levels used to initialize the $\mu_{k}$ in the adaptive noise schedule parameterization.
The experiments in Dieleman et al. (2022) show that $\sigma_{\text{init}}$ is a crucial hyperparameter for score interpolation on text data. The same sensitivity does not translate to the tabular data domain, as shown in our results in Table 7. The much smaller embedding dimension (16 vs. 256) and the feature-specific embeddings significantly decrease the number of distinguishable categories. Compared to a vocabulary size of 32000 for text data (Dieleman et al., 2022), we only face a maximum of 3151 categories in the lending dataset (see Table 3). Thus, unlike other generative (diffusion) models for tabular data, CDTD scales to a practically arbitrary number of categories.
Our proposed functional form for the adaptive noise schedules (see Appendix D) is the first to allow for the incorporation of prior information about the importance of low vs. high (normalized) noise levels. For this, we adjust the weight of low noise levels which directly determines the location of the inflection point $\mu_{k}$ (see Section 3.3). The results in Table 8 indicate low sample quality sensitivity to weight changes for a per-type noise schedule. The initialization only impacts the time to convergence but not (much) the location of the optimum. In our experiments, the number of training steps (30000) appears to be high enough for all model variants to converge.
Table 7: Performance sensitivity of CDTD (per type) to changes in the standard deviation $\sigma_{\text{init}}$ in the initialization of the embeddings of categorical features. Each metric is averaged over five seeds. As a robust measure, we report the median over the ablation study datasets acsincome, adult, beijing and churn .
$\sigma_{\text{init}}$ RMSE F1 AUC L2 distance of corr. Detection score JSD WD DCR
1 0.032 0.017 0.006 0.126 0.564 0.011 0.004 0.311
0.1 0.035 0.016 0.004 0.128 0.570 0.012 0.004 0.358
0.01 0.032 0.017 0.005 0.131 0.566 0.011 0.004 0.369
0.001 (default) 0.033 0.015 0.004 0.127 0.560 0.013 0.003 0.372
Table 8: Performance sensitivity of CDTD (per type) to changes in the prior weight of low noise levels in the initialization of the adaptive noise schedules. Each metric is averaged over five seeds. As a robust measure, we report the median over the ablation study datasets acsincome, adult, beijing and churn .
Weight RMSE F1 AUC L2 distance of corr. Detection score JSD WD DCR
1 0.036 0.015 0.004 0.143 0.651 0.015 0.003 0.313
2 0.030 0.014 0.005 0.147 0.651 0.014 0.003 0.352
3 (default) 0.033 0.015 0.004 0.154 0.651 0.013 0.004 0.366
4 0.034 0.019 0.005 0.148 0.656 0.013 0.004 0.370
Appendix M Advantages of Diffusion in Data Space
These days, inspired from diffusion models in the image and video domains, much work relies on the idea of latent diffusion. Here, we want to briefly discuss and emphasize that for tabular data, diffusion in latent space (represented by TabSyn) has important drawbacks and how CDTD, a diffusion model defined in data space differs from that.
Latent diffusion models first encode the data and map it into a latent space. The diffusion model itself is then trained in that latent space. Hence, the performance of the diffusion model directly depends on a second, separate model, with a separate training procedure. TabSyn uses a VAE model to encode mixed-type data into a common continuous space that is not lower-dimensional, so as to minimize reconstruction errors. Any reconstruction errors caused by the incapability of the VAE in turn reduce the sample quality of the eventually generated samples, no matter the capacity of the diffusion model. This suggests that we would want to train a high capable encoder/decoder, which adds additional training costs. Figure 3 shows that latent diffusion is not necessarily more efficient in the tabular data domain. In particular, if the latent space is not lower-dimensional to minimize reconstruction error, then sampling speed is not improved.
We further hypothesize that much tabular data, due to the lack of redundancy and spatial or sequential correlation, is difficult to summarize efficiently in a joint latent space. Hence, compared to other domains, larger VAEs and higher-dimensional latent spaces are required, increasing the training time. Also, there is the risk of the VAE not picking up on subtle correlations within the data or distorting existing correlations by mapping into the latent space. Any correlations not properly encoded in the latent space, cannot be learned or exploited by the diffusion model. Since we optimize the VAE on an average loss, its reconstruction and encoding performance of, for instance, minority classes or extreme values in long-tailed distributions is likely lacking. This makes the job of the diffusion model more difficult, if not impossible.
Lastly, we take great care in homogenizing categorical and continuous features throughout the training process (see Appendix A and B). This is a crucial part of modeling mixed-type data. Using a VAE to define a diffusion process in latent space only shifts the necessity for homogenization to the VAE training process. Not balancing different feature- or data-types and their losses induces an implicit importance weight for each feature. Thus, the VAE may sacrifice the reconstruction quality of some features in favor of others (Kendall et al., 2018; Ma et al., 2020).
To empirically investigate the difference of diffusion in data space (CDTD) and latent diffusion (TabSyn), we examine the worst feature-specific sample quality and other metrics that directly benefit from the model generating all features well. Our results in Table 9 show that, latent diffusion comes with a considerable decrease in sample quality (while imposing a similar architecture and number of parameters as well as sampling steps, see Appendix G). In particular, the attained maximum metrics indicate that TabSyn has issues modeling all features and their correlations sufficiently well. This supports our argument that a homogenization of data types is of crucial importance to avoid having the model implicitly favor one feature over another.
Table 9: A comparison of the CDTD model to latent diffusion (TabSyn). We average each metric over five sampling seeds and as a robust measure report the median over the ablation datasets acsincome, adult, beijing and churn. Abs. diff. in corr. matrices refers to the absolute differences in the correlation matrices between ground truth and synthetic data. The maximum, minimum and mean are taken across features.
Detection score L2 dist. of corr. JSD WD Abs. diff. in corr. matrices
min mean max min mean max min max
TabSyn 0.772 0.479 0.005 0.018 0.046 0.003 0.006 0.017 0 0.133
CDTD (per type) 0.566 0.131 0.001 0.012 0.022 0.001 0.003 0.007 0 0.052
improvement over TabSyn 1.364 3.656 5.000 1.500 2.091 3.000 2.000 2.429 0 2.558
Appendix N Comparison to Related Work
Table 10 summarizes our comparison of CDTD to the selected diffusion-based benchmark models, that is, TabSyn, TabDDPM and CoDi. Of those models, only TabSyn applies diffusion in latent space, which comes with both advantages and costs (as discussed in Appendix M). TabSyn is the only other model besides CDTD that avoids one-hot encoding categorical features by using embeddings. This improves the scalability to a higher number of categories without blowing up the input dimensions. Although both models utilize embeddings, TabSyn’s generative capabilities are more constrained by jointly encoding all features in a latent space. As such, it is still less flexible than CDTD, in particular when modeling very unbalanced categorical data. Information on rare categories may easily be cut off in favor of attributing more capacity in the latent space to more prominent categories or features. It should also be noted that TabSyn is the only model that makes use of a Transformer architecture in its VAE, which means that it scales quadratically in the number of features and therefore may not be easily scaled to high-dimensional data.
The CDTD model is the first to utilize adaptive and type- or feature-specific noise schedules to model tabular data. Further, we take great care in homogenizing categorical and continuous features throughout the training process, including the model initialization (see Appendix A and B). No other model attempts balancing the different features types. This is problematic as it suggests that other models may suffer from feature-specific induced implicit importance weights that impact both training and generation processes. Hence, the sample quality of some features may be unintentionally sacrificed in favor of increasing the sample quality of other features (Kendall et al., 2018; Ma et al., 2020). Note that this also applies to TabSyn: Even though their diffusion model avoids this issue by relying on a single type of loss due to the continuous latent space, the VAE training process does not account for any balancing issues between the two data types. Hence, the balancing issue is not eliminated but got only shifted to the encoder VAE.
Lastly, CDTD and TabSyn are the only models that define the diffusion process in continuous space. As such, other advanced techniques, like classifier-free guidance or ODE/SDE samplers, can be directly applied. To accommodate categorical data, CoDi and TabDDPM make use of multinomial diffusion (Hoogeboom et al., 2021), which is an inherently discrete process and therefore prohibits such applications.
Table 10: Comparison of CDTD to the diffusion-based generative models CoDi, TabDDPM and TabSyn. ($\ast$) Note that the VAE trained as part of the TabSyn model does not balance type-specific losses, which induces an implicit weighting among features. This can worsen the sample quality of some features in favor of others.
defined in feature space avoids one-hot encoding balances feature types adaptive noise schedule type- or feature- specific noise schedules diffusion in continuous space
CoDi
TabDDPM
TabSyn $\ast$
CDTD (ours)
Appendix O Examples of Learned Noise Schedules
@@
Next, we show the learned noise schedules for the smallest (churn) and the largest (acsincome) datasets. Additionally, we illustrate the fit of single, per type and per feature schedules to the respective losses.
@@
Appendix P Qualitative Comparisons
@@
Figure 10: Comparison of some univariate distributions for adult, bank, churn, nmes.
Figure 11: Comparison of some univariate distributions for beijing, default, lending, news. (Note that CoDi is prohibitively expensive to train on lending and therefore excluded.)
Appendix Q Visualizations of Captured Correlations
Figure 12: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the acsincome dataset. TabDDPM generates NaNs for this dataset and is therefore excluded. SMOTE takes too long for sampling. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
@@
Figure 17: Element-wise absolute differences of the correlation matrices between the real training set and the synthetic data for the covertype dataset. SMOTE takes too long for sampling. Continuous (cont.) and categorical (cat.) features are indicated on the axes.
@@
Appendix R Detailed Results